OpenStudy (anonymous):

how do you find the y intercept and the vertical asymptotes step by step of y=(4x)/(x^2-4) how do you find the y intercept and the vertical asymptotes step by step of y=(4x)/(x^2-4) @Mathematics

OpenStudy (amistre64):

the y intercept is always where the graph cross the y axis; and the value of x at the y axis is always 0

OpenStudy (amistre64):

set x=0 to find the value for y

OpenStudy (amistre64):

the verts are where the bottom goes bad; divide by zero is a vert

OpenStudy (anonymous):

ya but i need help with the algebra of actually solving the y int

OpenStudy (amistre64):

if you factor top and bottom and whats left over after like terms cancel defines your verts

OpenStudy (amistre64):

really? y=4x/(x^2-4) y=4.0/(0^2-4) y=0/-4 = 0

OpenStudy (anonymous):

isnt that the x intercept?

OpenStudy (amistre64):

.....no, that is the y intercept. It just so happens that it is ALSO the x intercept since it crosses at the origin

OpenStudy (amistre64):

but just set x=0 and solve for the y intercept; always.

OpenStudy (anonymous):

alright so how do you solve it setting y=0

OpenStudy (amistre64):

a fraction is only zero when the top of it is zero, right?

OpenStudy (amistre64):

so in this case when does 4x=0?

OpenStudy (anonymous):

ooooooooooooh alright thanks sorry I couldn't understand my teacher when she was explaining she made it sound a lot more complex

OpenStudy (amistre64):

job security ;)